OpenStudy (anonymous):

Find the length of the radius of a sphere with a surface area of 78.54 km2. A. 5 ft B. 15 ft C. 10 ft D. 2.5 ft

zepdrix (zepdrix):

The equation for Surface Area of a Sphere is given by:\[\huge A=4\pi r^2\] In this problem they gave us A, and we need to solve for r. \[\huge 78.54=4 \pi r^2\]Divide both sides by 4pi:\[\huge \frac{78.54}{4 \pi}=r^2\]Then, to solve for r, simply take the square root of both sides, yay!!

OpenStudy (anonymous):

how do you that?

zepdrix (zepdrix):

With a calculator of course!! ^^ If you don't have a Pi button on your calculator, then use 3.14 instead. Having trouble with which part? :O the square root step?